OpenStudy (anonymous):

The numbers 20,21,29 form a Pythagorean triple. Which other set of numbers is a Pythagorean triple? A.21,22,30 B.40,42,58 C.60,66,87 D.70,84,116

OpenStudy (anonymous):

you need to check if \[a^2+b^2=c^2\]

OpenStudy (anonymous):

how do i find the triple i dont get that part

OpenStudy (anonymous):

actually you don't. you can just double each entry and see that you get 40,42,58

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

pythagorean triple means 3 integers a, b, c, with \[a^2+b^2=c^2\]

OpenStudy (anonymous):

they are easy to find, i can show you how if you like